Calculus I, Fall 2013 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 13

1. Section 7.2: Problems 13, 19, 21.

2. Section 7.3: Problems 1, 11, 15, 19, 37, 45, 57.

3. Section 7.3: Derive Table 7.8 using formula for derivatives of inverse functions.